The current one-year Treasury rate is 5.23%, the two-year rate is 5.63%, and the three year rate is 4.84%. Assume expectations theory. What is the one-year rate expected two years from today (E(2r3)). Enter in percent form without the percent sign.
Answer #1

Based on Pure Expectations Theory,

(1 + 3 Yr Rate)3 = (1 + 2 Yr Rate)2 * (1 + 1 Yr Rate 2 Yr from now)

(1 + 4.84%)3 = (1 + 5.63%)2 * (1 +  1 Yr Rate 2 Yr from now)

1.152341 = 1.11577 * (1 +  1 Yr Rate 2 Yr from now)

(1 +  1 Yr Rate 2 Yr from now) = 1.0328

1 Yr Rate 2 Yr from now = 3.28%